With what tension must a rope with length 2.50m and mass 0.120 kg be stretched for transverse waves of frequency of 40.0 Hz to have a wavelength of 0.750m.
1
Expert's answer
2019-05-13T10:29:42-0400
"v=f\lambda=(40)(0.75)=30 \frac{m}{s}"

Also


"v=\sqrt{\frac{Tl}{m}}"

Tension:

"T=\frac{mv^2}{l}=\frac{0.12(30)^2}{2.5}=43.2\ N"